Accelerator Update - July 28-July 30
The integrated luminosity for last week was 9.7 pb-1.
MiniBooNE tripped off due to an airflow interlock at 11:35 AM. The experts investigating thought it was due to the effect of humidity on the electronics. The experiment was back on line an hour later.
At 8:36 PM, a Tevatron (TeV) sector A0 nitrogen Dewar had a relief valve open, to relieve pressure, but then reset slowly, which caused a small cloud of nitrogen to form. It had dissipated before a cryo system expert arrived and switched to another Dewar. The valve will be repaired later.
Operations terminated store 2826 at 9 PM.
Operators made the TeV ready for access at 10 PM. The Collider Detector Facility (CDF) and D-zero (D0) accessed their collision halls.
CDF came out of access at 11:18 PM.
Tuesday July 29
At 12:10 AM, techs reported finishing their work on the E2 Kautzky valve and F4 wet engine. At the same time vacuum techs were checking for vacuum leaks in the E2 system.
D0 came out of access at 2:42 AM.
Vacuum work in sector E2 finished at 4:07 AM.
Operations put the TeV through a dry squeeze at 5:42 AM. At 6:18 AM, Operations conducted a squeeze with beam.
Operations put the TeV into shot setup at 6:40 AM and at 8:48 AM established store 2828 with an initial luminosity of 3.7E31. Pbar resumed stacking.
The I- Source sparked at 9:54 AM.
Recycler experts took beam and stored it for studies.
At 10:54 AM, MiniBooNE tripped off due to a problem with a LCW system. Experts investigated. They had the system reset by a little after 11 AM.
A Switchyard expert began a study period at 8:54 PM.
A TeV RF station (TRF7) tripped off at 9:37 PM. Operators responded to the F0 service building to investigate, but first had to distract a Fox snake before getting inside. Eventually, the station was reset.
(A fox snake can grow to from 36 to 56 inches long. It is a non-poisonous snake that kills rodents by constriction. The younger snakes may coil and strike when cornered, and may vibrate its tail, though it has no real rattle. This, unfortunately, often gets the snake killed by those who can't tell the difference between it and a rattlesnake.)
Wednesday July 30
A bad CAMAC card at 12:52 AM halted beam to the Meson area until it was replaced.
At 2:49 AM, Operations could not extract beam from the Pbar Accumulator. While operators investigated, stacking resumed normal operations ten minutes later. Then, at 3:15 AM, it happened again.
A Booster power supply (West Anode) and the Tevatron Electron Lens (TEL) tripped off at 3:20 AM. A MI RF station (MIRF5) also tripped off. Everything reset eventually.
Operations terminated store 2828 at 4:09 AM and then put the TeV into a dry squeeze.
A Booster RF station (BRF11) tripped off at 4:36 AM and cause another power supply (WAPS) to trip off also. Operators bypassed BRF11 and turn the Booster back on at 4:51 AM, but WAPS tripped off again a few minutes later. Operators called in an expert. He said that he could not fix it before the next shot, so Operations will temporarily run Booster with out it.
Cryo techs at CHL told Operations at 5:38 AM that the A0 cryo compressor had tripped off. The cryo coordinator advised them to turn on another compressor.
Operation established store 2830 at 7:30 AM with an initial luminosity of 3.5E31.
Comments and Suggestions
|last modified 7/30/2003 email Fermilab|